OpenStudy (anonymous):

[ 0] [ 0 ] [ 4] v1=[ 0] v2=[-3] v3=[-2] [-3] [ 9 ] [-6] Does {v1,v2,v3} span R^3? Why or why not?

OpenStudy (anonymous):

what do you mean by that? does it mean if it has a solution or not?

OpenStudy (unklerhaukus):

\[v_1=\left[\begin{array}\ 0\\0\\-3\end{array}\right]\qquad\qquad v_2=\left[\begin{array}\ 0\\-3\\9\end{array}\right]\qquad\qquad v_3=\left[\begin{array}\ 4\\-2\\-6\end{array}\right]\]

OpenStudy (anonymous):

im not sure what it means

OpenStudy (anonymous):

i think its asking if they are linearly independent or not